# SU(3) gauge field of magnons in antiferromagnetic skyrmion crystals

### v2

@@ Abstract
Quasiparticle excitations in material solids often experience a fictitious gauge field, which can be a potential source of intriguing transport phenomena. Here, we show that low-energy excitations in insulating antiferromagnetic skyrmion crystals on the triangular lattice are effectively described by magnons with an SU(3) gauge field. The three-sublattice structure in the antiferromagnetic skyrmion crystals is inherited as three internal degrees of freedom for the magnons, which are coupled with their kinetic motion via the SU(3) gauge field that arises from the topologically nontrivial spin texture in real space. We also demonstrate that the non-commutativity of the SU(3) gauge field breaks an effective time-reversal symmetry and contributes to a magnon thermal Hall effect.
@@
The DM interactions or noncoplanar spin textures typically yield a zero net flux. In this case, the magnon systems adhere to the no-go condition that precludes the magnon thermal Hall effect in edge-shared lattice geometries such as the square and triangular lattices [12, 18]. There, due to the geometrically equivalent cells with opposite fluxes in nearest neighbors, the systems have the effective time-reversal symmetry that leaves the flux pattern unchanged while converting the sign of the thermal Hall conductivity [12, 18]. In contrast, lattices that feature corner-sharing, such as the kagome and pyrochlore lattices, escape this scenario; their geometrically inequivalent neighboring cells allow for the finite thermal Hall conductivity [12, 18].
@@
Ferromagnetic skyrmion crystals (FM-SkXs), characterized by their topologically nontrivial swirling spin textures, yield a finite net flux, thereby contributing to the anomalous transports even in the edge-shared lattices [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. The magnon thermal Hall effect is observed in the FM-SkX phase of GaV4Se8 [29], where (V4Se4)5+ clusters form the triangular-lattice FM-SkX in the $[111]$ plane [30, 31, 32, 33]. Recently, the magnon thermal Hall effect is also observed in the antiferromagnetic skyrmion crystal (AFM-SkX) phase of MnSc2S4 [34], with Mn2+ ions forming the triangular-lattice AFM-SkX in the $[111]$ plane [35, 36, 37, 38]. The AFM-SkXs on the triangular lattice consist of three intertwined FM-SkXs that are antiferromagnetically coupled, leading to a three-sublattice structure [39, 40, 41, 42, 43]. The net flux in the AFM-SkXs, however, is naively expected to be zero because of their antiferromagnetic coupling. To explain the origin of the magnon thermal Hall effect in the AFM-SkXs, the authors in Ref. [34] introduce the concept of an SU(3) gauge field of magnons. However, how the spin textures in the AFM-SkXs are translated into the SU(3) gauge field has not yet been clarified well.
Figure 1: Skyrmion crystals and gauge fields of magnons. (a), (b) Schematic illustration of the real-space spin configuration of the (a) FM-SkXs and (b) AFM-SkXs on the triangular lattice. The AFM-SkXs consist of three intertwined FM-SkXs as shown in different colors, giving rise to the three-sublattice structure. (c), (d) Schematic illustration of the uniform flux in the low-energy magnon systems for the (c) FM-SkXs and (d) AFM-SkXs. The clockwise (counterclockwise) arrows indicate the negative (positive) direction of the flux. The three different colors on each site in (d) represent the three internal degrees of freedom for the magnons. The uniform flux $\Phi$ originates from the commutation relation of an SU(3) gauge field.
@@
In this Letter, we bridge this gap by constructing the effective field theory of magnons in the AFM-SkXs from a spin model on the triangular lattice. We show that the three sublattices in the AFM-SkXs introduce the three internal degrees of freedom for the magnons, which are coupled with their kinetic motion via the SU(3) gauge field originating from the combined effect of the spin textures of three FM-SkXs and the local antiferromagnetic coupling between them. We also find that the commutation relation of the SU(3) gauge field generates a uniformly distributed flux, which is responsible for the magnon thermal Hall effect.
@@
Effective field theory of magnons in FM-SkXs.—Before we explore the effective field theory of magnons in the AFM-SkXs, it is instructive first to briefly discuss that in the FM-SkXs [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. We start from a spin Hamiltonian on the triangular lattice,
^$\hat{\mathcal{H}}=J\sum_{\braket{i,j}}\hat{\bm{S}}_{i}\cdot\hat{\bm{S}}_{j}+\cdots,$ (1)
where $\braket{i,j}$ denotes the pair of nearest-neighbor $i$ and $j$ sites, $\hat{\bm{S}}_{i}=(\hat{S}_{i}^{x},\hat{S}_{i}^{y},\hat{S}_{i}^{z})$ is the spin-$S$ operator at site $i$, $J$ is the Heisenberg exchange coupling constant, and ellipsis indicates terms that stabilize the FM-SkXs (AFM-SkXs) for $J<0$ ($J>0$) such as a DM interaction, magnetic field, or single-ion anisotropy. We only focus on the Heisenberg exchange interaction term since the other terms do not essentially affect the following discussion.
@@
Given the slow spatial variation of spin orientations in the FM-SkXs (see Fig. 1(a)), the low-energy theory is expected to be described by a continuously varying spin-density operator $\hat{\bm{s}}(\bm{r})$. Adopting this assumption, the effective field theory is derived by substituting $\hat{\bm{S}}_{i}$ with $v\hat{\bm{s}}(\bm{r})$ and $\sum_{i}$ with $(1/v)\int d^{2}\bm{r}$, where $v=\sqrt{3}a^{2}/2$ is the volume per site with lattice constant $a$. Applying these substitutions to Eq. (1) leads to the effective Hamiltonian
^$\hat{\mathcal{H}}_{\mathrm{eff}}^{\mathrm{FM}}=3Jv\int d^{2}\bm{r}\ \hat{\bm{s}}(\bm{r})\cdot\left(1+\frac{a^{2}}{4}\nabla^{2}\right)\hat{\bm{s}}(\bm{r}),$ (2)
@@
where we drop higher-order derivative terms, which are irrelevant at low energies. To describe the magnon excitations, we employ the Holstein-Primakoff transformation [44]
@@
where $\hat{b}(\bm{r})$ ($\hat{b}^{\dagger}(\bm{r})$) is the magnon annihilation (creation) operator at $\bm{r}$, $\bm{m}(\bm{r})$ is the unit vector pointing in the direction of the spin in the classical ground state, and $\bm{e}^{\pm}(\bm{r})=(\bm{e}^{x}(\bm{r})\pm i\bm{e}^{y}(\bm{r}))/\sqrt{2}$ with the unit vectors, $\bm{e}^{x}(\bm{r})$ and $\bm{e}^{y}(\bm{r})$, satisfying $\bm{e}^{x}(\bm{r})\times\bm{e}^{y}(\bm{r})=\bm{m}(\bm{r})$. The three unit vectors, $\bm{e}^{x}(\bm{r})$, $\bm{e}^{y}(\bm{r})$, and $\bm{m}(\bm{r})$, form a local orthonormal basis (see Fig. 2(a)). We neglect constant terms, terms quadratic in the derivative of $\bm{e}^{\pm}(\bm{r})$, and magnon-magnon interaction terms, which are only relevant at high energies [45, 46]. From Eqs. (2) and (3), we obtain the effective magnon Hamiltonian as
@@
^$\hat{\mathcal{H}}_{\mathrm{eff}}^{\mathrm{FM}}\simeq\frac{3}{2}JSa^{2}\int d^{2}\bm{r}\ \hat{b}^{\dagger}(\bm{r})\left(\bm{\nabla}-i\bm{A}(\bm{r})\right)^{2}\hat{b}(\bm{r}),$ (4)
@@
with a U(1) gauge field $A_{\mu}(\bm{r})=i\bm{e}^{+}(\bm{r})\cdot\partial_{\mu}\bm{e}^{-}(\bm{r})$ ($\mu=x,y$). The associated fictitious magnetic field, $B(\bm{r})=\partial_{x}A_{y}(\bm{r})-\partial_{y}A_{x}(\bm{r})$, is calculated as [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]
$B(\bm{r})=\bm{m}(\bm{r})\cdot\left(\partial_{x}\bm{m}(\bm{r})\times\partial_{y}\bm{m}(\bm{r})\right),$ (5)
which takes the finite value due to the spatial variations in $\bm{m}(\bm{r})$. Therefore, the low-energy excitations in the FM-SkXs are described by the magnons coupled with the U(1) gauge field, which arises from the spatially varying spin texture in real space [20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
@@
Importantly, the topologically nontrivial spin texture in the FM-SkXs induces a uniform component of the fictitious magnetic field, $\bar{B}=(1/V)\int d^{2}\bm{r}\ B(\bm{r})$, which is proportional to the skyrmion density $(1/4\pi V)\int d^{2}\bm{r}\ \bm{m}(\bm{r})\cdot(\partial_{x}\bm{m}(\bm{r})\times\partial_{y}\bm{m}(\bm{r}))$, where $V$ is the system’s volume. The finite $\bar{B}$ generates magnon Landau levels and contributes to the magnon thermal Hall effect [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. In terms of the original lattice, the finite $\bar{B}$ gives rise to the uniformly distributed flux $\phi=\bar{B}v/2$ as illustrated in Fig. 1(c), which breaks the effective time-reversal symmetry that prohibits the finite thermal Hall conductivity [12, 18].
We also comment on gauge redundancy. The presence of $\bm{A}(\bm{r})$ in Eq. (4) does not necessarily indicate the emergence of the U(1) gauge field, as it can also be artificially introduced by a local U(1) gauge transformation, $\hat{b}(\bm{r})\to\mathrm{e}^{i\varphi(\bm{r})}\hat{b}(\bm{r})$ and $\bm{A}(\bm{r})\to\bm{A}(\bm{r})+\bm{\nabla}\varphi(\bm{r})$. The latter transformation is equivalent to the rotation of the local orthonormal basis by $-\varphi(\bm{r})$ around $\bm{m}(\bm{r})$, leading to $\bm{e}^{\pm}(\bm{r})\to\mathrm{e}^{\pm i\varphi(\bm{r})}\bm{e}^{\pm}(\bm{r})$. The local U(1) gauge transformation originates from the redundancy of the description and hence all physical quantities should be gauge invariant. Therefore, the emergence of the U(1) gauge field should be characterized by finite $B(\bm{r})$, not $\bm{A}(\bm{r})$.
@@
Figure 2: Local orthonormal basis. (a) Local orthonormal basis in the FM-SkXs. The spatially varying $\bm{e}^{x}(\bm{r})$ and $\bm{e}^{y}(\bm{r})$ lead to the finite $\bm{A}(\bm{r})$ in the effective magnon Hamiltonian (4). (b) Three local orthonormal bases in the AFM-SkXs. The three unit vectors $\bm{m}_{\ell}(\bm{r})$ ($\ell=1,2,3$) are on the same plane due to the local constraint $\sum_{\ell}\bm{m}_{\ell}(\bm{r})=\bm{0}$. We take the unit vector $\bm{n}(\bm{r})=\bm{e}_{\ell}^{x}(\bm{r})$ to be orthogonal to this plane.
@@
Effective field theory of magnons in AFM-SkXs.—We turn to the effective field theory of magnons in the AFM-SkXs. A key observation is that, since the AFM-SkXs consist of three intertwined FM-SkXs, we need to introduce three continuously varying spin-density operators $\hat{\bm{s}}_{\ell}(\bm{r})$ with sublattice index $\ell=1,2,3$ to describe the low-energy excitations [34]. The spin operator $\hat{\bm{S}}_{i}$ is substituted with $v^{\prime}\hat{\bm{s}}_{\ell}(\bm{r})$ when site $i$ belongs to the sublattice $\ell$, and $\sum_{i}$ is substituted with $(1/v^{\prime})\int d^{2}\bm{r}\sum_{\ell}$, where $v^{\prime}=3v$. By these substitutions, we obtain the effective Hamiltonian from Eq. (1) as
^$\hat{\mathcal{H}}_{\mathrm{eff}}^{\mathrm{AFM}}=\frac{3Jv^{\prime}}{2}\int d^{2}\bm{r}\sum_{\ell}\sum_{\ell^{\prime}\neq\ell}\hat{\bm{s}}_{\ell}(\bm{r})\cdot\left(1+\frac{a^{2}}{4}\nabla^{2}\right)\hat{\bm{s}}_{\ell^{\prime}}(\bm{r}).$ (6)
@@
As in the case of the FM-SkXs, we define a local orthonormal basis, $\bm{e}_{\ell}^{x}(\bm{r})$, $\bm{e}_{\ell}^{y}(\bm{r})$, and $\bm{m}_{\ell}(\bm{r})$, with sublattice index $\ell$ as shown in Fig. 2(b). The unit vector $\bm{m}_{\ell}(\bm{r})$ denotes the direction of the spin at $\bm{r}$ with index $\ell$. To incorporate the local $120^{\circ}$ order in the AFM-SkXs, we impose the local constraint $\sum_{\ell}\bm{m}_{\ell}(\bm{r})=\bm{0}$, indicating that the three unit vectors $\bm{m}_{\ell}(\bm{r})$ are on the same plane. We introduce the unit vector $\bm{n}(\bm{r})$ that is orthogonal to this plane as $\bm{n}(\bm{r})=(2/\sqrt{3})\bm{m}_{1}(\bm{r})\times\bm{m}_{2}(\bm{r})$, and take $\bm{e}_{\ell}^{x}(\bm{r})=\bm{n}(\bm{r})$ for all $\ell=1,2,3$, which simplifies the following calculations. Applying the Holstein-Primaloff transformation (3) with index $\ell$ and unit volume $v^{\prime}$ to Eq. (6) leads to the effective magnon Hamiltonian
@@
^$\hat{\mathcal{H}}_{\mathrm{eff}}^{\mathrm{AFM}}\simeq\frac{1}{2}\int d^{2}\bm{r}\sum_{\ell,\ell^{\prime}}\begin{pmatrix}\hat{b}_{\ell}^{\dagger}(\bm{r})&\hat{b}_{\ell}(\bm{r})\end{pmatrix}\tilde{H}_{\ell,\ell^{\prime}}(\bm{r})\begin{pmatrix}\hat{b}_{\ell^{\prime}}(\bm{r})\\ \hat{b}_{\ell^{\prime}}^{\dagger}(\bm{r})\end{pmatrix},$ (7)
@@
where $\hat{b}_{\ell}(\bm{r})$ ($\hat{b}_{\ell}^{\dagger}(\bm{r})$) is the magnon annihilation (creation) operator with sublattice index $\ell$ and $\tilde{H}_{\ell,\ell^{\prime}}(\bm{r})$ is the $2\times 2$ matrix defined as $\tilde{H}_{\ell,\ell}(\bm{r})=3JS\tau^{0}$ for $\ell=\ell^{\prime}$ and

### v3

# Thermal Hall effect and gauge-field picture of magnons in antiferromagnetic skyrmion crystals

@@ Abstract
Quasiparticle excitations in material solids often experience a fictitious gauge field, which can be a potential source of intriguing transport phenomena. Here, we show that low-energy excitations in insulating antiferromagnetic skyrmion crystals on the triangular lattice are effectively described by magnons with an SU(3) gauge field. The three-sublattice structure in the antiferromagnetic skyrmion crystals is inherited as three internal degrees of freedom for the magnons, which are coupled to their kinetic motion via the SU(3) gauge field that arises from the topologically nontrivial spin texture in real space. We show that the non-commutativity of the SU(3) gauge field breaks an effective time-reversal symmetry and contributes to a magnon thermal Hall effect. We further demonstrate the emergence of the finite thermal Hall conductivity in the antiferromagnetic skyrmion crystals by the linear spin-wave theory. The possible impact of different gauge structures on a thermal Hall conductivity is also discussed.
@@
The DM interactions or noncoplanar spin textures typically yield a staggered flux pattern with zero net flux. In this case, the magnon systems adhere to the no-go condition that precludes the magnon thermal Hall effect in edge-shared lattice geometries such as the square and triangular lattices [12, 18]. There, due to the geometrically equivalent cells with opposite fluxes in nearest neighbors, the systems have the effective time-reversal symmetry that leaves the flux pattern unchanged while converting the sign of the thermal Hall conductivity [12, 18]. In contrast, lattices that feature corner-sharing, such as the kagome and pyrochlore lattices, escape this scenario; their geometrically inequivalent neighboring cells allow for the finite thermal Hall conductivity [12, 18].
@@
Ferromagnetic skyrmion crystals (FM-SkXs), characterized by their topologically nontrivial swirling spin textures, yield a finite net flux, thereby contributing to the anomalous transports even in the edge-shared lattices [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. The magnon thermal Hall effect is observed in the FM-SkX phase of GaV4Se8 [29], where (V4Se4)5+ clusters form the triangular-lattice FM-SkX in the $[111]$ plane [30, 31, 32, 33]. Recently, the magnon thermal Hall effect has also been observed in the antiferromagnetic skyrmion crystal (AFM-SkX) phase of MnSc2S4 [34], with Mn2+ ions forming the triangular-lattice AFM-SkX in the $[111]$ plane [35, 36, 37, 38]. The AFM-SkXs on the triangular lattice consist of three intertwined FM-SkXs that are antiferromagnetically coupled, leading to a three-sublattice structure [39, 40, 41, 42, 43]. The net flux in the AFM-SkXs, however, is naively expected to be zero because of their antiferromagnetic coupling. To explain the origin of the magnon thermal Hall effect in the AFM-SkXs, the authors in Ref. [34] introduce the concept of an SU(3) gauge field of magnons. However, how the spin textures in the AFM-SkXs are translated into the SU(3) gauge field has not yet been clarified well.
@@
In this Letter, we bridge this gap by constructing the effective field theory of magnons in the AFM-SkXs from a spin model on the triangular lattice. We show that the three sublattices in the AFM-SkXs introduce the three internal degrees of freedom for the magnons, which are coupled to their kinetic motion via the SU(3) gauge field originating from the combined effect of the spin textures of three FM-SkXs and the local antiferromagnetic coupling between them. We find that the commutation relation of the SU(3) gauge field generates a uniformly distributed flux, which is responsible for the magnon thermal Hall effect. Using the linear spin-wave theory, we also demonstrate that the thermal Hall conductivity remains very small in the helical phase but becomes substantially large in the AFM-SkX phase. The possible impact of the different gauge structures in FM-SkXs and AFM-SkXs on the thermal Hall conductivity is also discussed.
@@
Effective field theory of magnons in FM-SkXs.—Before we explore the effective field theory of magnons in the AFM-SkXs, it is instructive first to briefly review that in the FM-SkXs [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. We start from a spin Hamiltonian on the triangular lattice,
$\hat{\mathcal{H}}=\sum_{\braket{i,j}}\left[J\hat{\bm{S}}_{i}\cdot\hat{\bm{S}}_{j}+\bm{D}_{i,j}\cdot(\hat{\bm{S}}_{i}\times\hat{\bm{S}}_{j})\right]-g\mu_{\mathrm{B}}B\sum_{i}\hat{S}_{i}^{z},$ (1)
where $\braket{i,j}$ denotes the pair of nearest-neighbor $i$ and $j$ sites, $\hat{\bm{S}}_{i}=(\hat{S}_{i}^{x},\hat{S}_{i}^{y},\hat{S}_{i}^{z})$ is the spin-$S$ operator at site $i$, $J$ is the Heisenberg exchange coupling constant, $\bm{D}_{i,j}=D\bm{e}^{Z}\times\bm{e}_{i,j}$ is the DM vector, $\bm{e}^{Z}$ is the unit vector pointing in the $z$-direction in spin space, $\bm{e}_{i,j}$ is the unit vector pointing from site $i$ to site $j$, and $B$ is the magnetic field in $z$-direction with g-factor $g$ and Bohr magneton $\mu_{\mathrm{B}}$. Various spin configurations can be realized in noncentrosymmetric magnets by the competition between exchange parameters, anisotropy, and magnetic field [44, 45]. For sufficiently large $D/|J|$ and $g\mu_{\mathrm{B}}B/|J|$, the FM-SkX (AFM-SkX) phase is realized in the ground state of Eq. (1) for $J<0$ (J>0J>0[39, 40, 46]. The schematic illustrations of the FM-SkX and AFM-SkX are shown in Fig. 2(a) and Fig. 2(b).
@@
Given the slow spatial variation of spin orientations in the FM-SkXs (see Fig. 2(a)), the low-energy theory is expected to be described by a continuously varying spin-density operator $\hat{\bm{s}}(\bm{r})$. Adopting this assumption, the effective field theory is derived by substituting $\hat{\bm{S}}_{i}$ with $v\hat{\bm{s}}(\bm{r})$ and $\sum_{i}$ with $(1/v)\int\mathrm{d}^{2}\bm{r}$, where $v=\sqrt{3}a^{2}/2$ is the volume per site with lattice constant $a$. Applying these substitutions to Eq. (1) leads to the effective Hamiltonian
$\hat{\mathcal{H}}_{\mathrm{eff}}^{\mathrm{FM}}=\int\mathrm{d}^{2}\bm{r}\left[\hat{\bm{s}}^{T}(\bm{r})K(\bm{r})\hat{\bm{s}}(\bm{r})-g\mu_{\mathrm{B}}B\hat{s}^{z}(\bm{r})\right],$ (2)
where $K(\bm{r})$ is the $3\times 3$ matrix defined as
$K(\bm{r})=3Jv\left(1+\frac{a^{2}}{4}\nabla^{2}\right)I+\frac{3}{2}Dva\begin{pmatrix}0&0&-\partial_{x}\\ 0&0&-\partial_{y}\\ \partial_{x}&\partial_{y}&0\end{pmatrix}$, (3)
@@
with the $3\times 3$ identity matrix $I$. Here, we drop higher-order derivative terms, which are irrelevant at low energies. To describe the magnon excitations, we employ the Holstein-Primakoff transformation [47]
@@
where $\hat{b}(\bm{r})$ ($\hat{b}^{\dagger}(\bm{r})$) is the magnon annihilation (creation) operator at $\bm{r}$, $\bm{m}(\bm{r})=(m^{x}(\bm{r}),m^{y}(\bm{r}),m^{z}(\bm{r}))$ is the unit vector pointing in the direction of the spin in the classical ground state, and $\bm{e}^{\pm}(\bm{r})=(\bm{e}^{x}(\bm{r})\pm i\bm{e}^{y}(\bm{r}))/\sqrt{2}$ with the unit vectors, $\bm{e}^{x}(\bm{r})$ and $\bm{e}^{y}(\bm{r})$, satisfying $\bm{e}^{x}(\bm{r})\times\bm{e}^{y}(\bm{r})=\bm{m}(\bm{r})$. The three unit vectors, $\bm{e}^{x}(\bm{r})$, $\bm{e}^{y}(\bm{r})$, and $\bm{m}(\bm{r})$, form a local orthonormal basis. We neglect constant terms, terms quadratic in the derivative of $\bm{e}^{\pm}(\bm{r})$, and magnon-magnon interaction terms, which are only relevant at high energies [48, 49]. From Eqs. (2) and (4), we obtain the effective magnon Hamiltonian as
@@
$\hat{\mathcal{H}}_{\mathrm{eff}}^{\mathrm{FM}}\simeq\int\mathrm{d}^{2}\bm{r}\ \hat{b}^{\dagger}(\bm{r})\left[\frac{3}{2}JSa^{2}\left(\bm{\nabla}-i\bm{A}(\bm{r})\right)^{2}-\hbar A_{0}(\bm{r})\right]\hat{b}(\bm{r}),$ (5)
@@
with U(1) gauge fields $A_{0}(\bm{r})=-g\mu_{\mathrm{B}}Bm^{z}(\bm{r})/\hbar$, $A_{x}(\bm{r})=i\bm{e}^{+}(\bm{r})\cdot\partial_{x}\bm{e}^{-}(\bm{r})+(D/Ja)m^{y}(\bm{r})$, and $A_{y}(\bm{r})=i\bm{e}^{+}(\bm{r})\cdot\partial_{y}\bm{e}^{-}(\bm{r})-(D/Ja)m^{x}(\bm{r})$. The Mermin-Ho relation [50] expresses the associated fictitious magnetic field, $B(\bm{r})=\partial_{x}A_{y}(\bm{r})-\partial_{y}A_{x}(\bm{r})$, in terms of $\bm{m}(\bm{r})$ as
$B(\bm{r})=\bm{m}(\bm{r})\cdot\left(\partial_{x}\bm{m}(\bm{r})\times\partial_{y}\bm{m}(\bm{r})\right)-\frac{D}{Ja}\sum_{\mu=x,y}\partial_{\mu}m^{\mu}(\bm{r}),$ (6)
@@
which is the gauge-invariant, physical field coupled to the magnons [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Its uniform component is proportional to the skyrmion density $\rho_{\mathrm{FM-SkX}}=(1/4\pi V)\int\mathrm{d}^{2}\bm{r}\ \bm{m}(\bm{r})\cdot(\partial_{x}\bm{m}(\bm{r})\times\partial_{y}\bm{m}(\bm{r}))$ and takes finite value in the FM-SkX, where $V$ is the system’s volume. In terms of the original lattice, the uniform component gives rise to uniformly distributed flux $\phi=2\pi v\rho_{\mathrm{FM-SkX}}$ as illustrated in Fig. 2(c). Therefore, the situation is analogous to that of charged particles moving in a uniform magnetic field, and the magnon thermal Hall effect emerges in a manner similar to the well-known electronic Hall effect [20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
The emergence of the uniform flux is particularly important since otherwise the no-go condition precludes a finite thermal Hall conductivity [12, 18]. We briefly review this fact by considering the triangular-lattice ferromagnet with the DM interaction shown in Fig. 2(e). The DM interaction leads to the complex hopping of magnons $\mathrm{e}^{i\phi_{\mathrm{D}}/3}\hat{b}_{i}^{\dagger}\hat{b}_{j}$ with $\phi_{\mathrm{D}}/3=-\arctan(D/J)$, resulting in the staggered flux pattern shown in Fig. 2(f). The flux pattern is preserved under a $\pi$-rotation along any straight line in the triangular lattice. This rotation, however, converts the sign of the thermal Hall conductivity, indicating that the DM-induced flux does not contribute to the magnon thermal Hall effect.
@@
Figure 1: Model and orthogonal basis. (a) Schematic illustration of the spin model on the triangular lattice. The DM vectors $\bm{D}_{i,j}$ are marked by blue arrows and the corresponding bond directions $\bm{e}_{i,j}$ by black arrows. (b) Three local orthonormal bases in the AFM-SkXs. The three unit vectors $\bm{m}_{\ell}(\bm{r})$ ($\ell=1,2,3$) are approximately on the same plane. We take the unit vector $\bm{n}(\bm{r})=\bm{e}_{\ell}^{x}(\bm{r})$ to be orthogonal to this plane.
@@
Effective field theory of magnons in AFM-SkXs.—We turn to the effective field theory of magnons in the AFM-SkXs. A key observation is that, since the AFM-SkXs consist of three intertwined FM-SkXs, we need to introduce three continuously varying spin-density operators $\hat{\bm{s}}_{\ell}(\bm{r})$ with sublattice index $\ell=1,2,3$ to describe the low-energy excitations [34]. The spin operator $\hat{\bm{S}}_{i}$ is substituted with $v^{\prime}\hat{\bm{s}}_{\ell}(\bm{r})$ when site $i$ belongs to the sublattice $\ell$, and $\sum_{i}$ is substituted with $(1/v^{\prime})\int\mathrm{d}^{2}\bm{r}\sum_{\ell}$, where $v^{\prime}=3v$. By these substitutions, we obtain the effective Hamiltonian from Eq. (1) as
$\hat{\mathcal{H}}_{\mathrm{eff}}^{\mathrm{AFM}}=\hskip-3.0pt\int\mathrm{d}^{2}\bm{r}\hskip-1.0pt\sum_{\ell}\left[\frac{3Jv^{\prime}}{2}\hskip-4.0pt\sum_{\ell^{\prime}(\neq\ell)}\hat{\bm{s}}_{\ell}^{T}(\bm{r})K(\bm{r})\hat{\bm{s}}_{\ell^{\prime}}(\bm{r})-g\mu_{\mathrm{B}}B\hat{s}_{\ell}^{z}(\bm{r})\right].$ (7)
@@
As in the case of the FM-SkXs, we define a local orthonormal basis, $\bm{e}_{\ell}^{x}(\bm{r})$, $\bm{e}_{\ell}^{y}(\bm{r})$, and $\bm{m}_{\ell}(\bm{r})$, with sublattice index $\ell$ as shown in Fig. 1(b). The unit vector $\bm{m}_{\ell}(\bm{r})$ denotes the direction of the spin at $\bm{r}$ with index $\ell$. To incorporate the local $120^{\circ}$ order in the AFM-SkXs, we impose the local constraint $\sum_{\ell}\bm{m}_{\ell}(\bm{r})\simeq 3m_{\mathrm{u}}^{z}\bm{e}^{Z}$ with a small uniform magnetization $m_{\mathrm{u}}^{z}\ll 1$, indicating that the three unit vectors $\bm{m}_{\ell}(\bm{r})$ are approximately on the same plane. We introduce the unit vector $\bm{n}(\bm{r})$ that is orthogonal to this plane as $\bm{n}(\bm{r})=(2/\sqrt{3})\bm{m}_{1}(\bm{r})\times\bm{m}_{2}(\bm{r})$, and take $\bm{e}_{\ell}^{x}(\bm{r})=\bm{n}(\bm{r})$ for all $\ell=1,2,3$, which simplifies the following calculations. Applying the Holstein-Primaloff transformation (4) with index $\ell$ and unit volume $v^{\prime}$ to Eq. (7) leads to the effective magnon Hamiltonian
@@
$\hat{\mathcal{H}}_{\mathrm{eff}}^{\mathrm{AFM}}\simeq\frac{1}{2}\int\mathrm{d}^{2}\bm{r}\sum_{\ell,\ell^{\prime}}\begin{pmatrix}\hat{b}_{\ell}^{\dagger}(\bm{r})&\hat{b}_{\ell}(\bm{r})\end{pmatrix}\tilde{H}_{\ell,\ell^{\prime}}(\bm{r})\begin{pmatrix}\hat{b}_{\ell^{\prime}}(\bm{r})\\ \hat{b}_{\ell^{\prime}}^{\dagger}(\bm{r})\end{pmatrix}$, (8)
@@
where $\hat{b}_{\ell}(\bm{r})$ ($\hat{b}_{\ell}^{\dagger}(\bm{r})$) is the magnon annihilation (creation) operator with sublattice index $\ell$ and $\tilde{H}_{\ell,\ell^{\prime}}(\bm{r})$ is the $2\times 2$ matrix defined as $\tilde{H}_{\ell,\ell}(\bm{r})=[3JS+g\mu_{\mathrm{B}}Bm_{\ell}^{z}(\bm{r})]\tau^{0}$ for $\ell=\ell^{\prime}$ and